# Adaptive Gradient Normalization and Independent Sampling for (Stochastic) Generalized-Smooth Optimization

### v2

@@ Appendix A Ablation Study
In order to have a comprehensive understanding of the performance of IAN-SGD in practical problems, we conduct ablation study on Phase Retrieval and DRO regarding important components of IAN-SGD, including adaptive normalization, batch size of independent samples and numerical stabilizer $\delta$.
@@ A.1 Effects of $\beta$
In the first experiment, we choose a unified normalization parameter $\beta=\frac{2}{3}$ for all the normalized methods, i.e., NSGD, NSGD with momentum, clipped SGD, SPIDER and IAN-SGD. To guarantee convergence, we adjust the learning rate accordingly, i.e., $\gamma=0.03$ for NSGD and NSGD with momentum, $\gamma=0.05$ for SPIDER, $\gamma=0.17$ for both clipped SGD and IAN-SGD. To make a fair comparison, we keep other parameters unchanged. Figure 4 (left) shows the comparison of objective value versus sample complexity for the Phase Retrieval problem. It can be observed that, by adjusting $\beta=\frac{2}{3}$, the objective value optimized by all algorithms decreases much faster compared with Figure 2. This indicates that adaptive normalization can accelerate convergence. Moreover, compared with other normalization methods, although IAN-SGD requires additional sampling at each iteration, the training loss still decreases faster than NSGD SGD with momentum and SPIDER.
Similarly, for DRO, we unify the normalization parameter of all the normalized methods, i.e., NSGD, NSGD with momentum, clipped SGD, SPIDER and IAN-SGD with $\beta=\frac{2}{3}$. To guarantee algorithm convergence, we adjust learning rate correspondingly, for NSGD, NSGDm and SPIDER, we keep learning rate unchanged, for clipped SGD, and IAN-SGD, we set $\gamma=0.08$. To make a fair comparison, we keep other parameters unchanged. Figure 4 (right) shows the comparison of objective value versus sample complexity. It can be observed that, by setting $\beta=\frac{2}{3}$, the objective value optimized by all normalization methods decreases faster than Figure 2. This verifies the effectiveness of adaptive normalization. Although IAN-SGD requires additional sampling, it converges faster than other normalized methods.
@@
In the second experiment, we vary $\beta$ for IAN-SGD and keep other parameters unchanged. Figure 5 shows the convergence result with different $\beta$ for Phase Retrieval and DRO. It can be observed that decreasing $\beta$ accelerates convergence in general. However, small $\beta$ can make convergence unstable. In the DRO experiment shown in Figure 5 (right), we observed that when $\beta=\frac{3}{5}$, the objective value curve fluctuates a lot. If $\beta$ is smaller than $\frac{3}{5}$, the algorithm fails to converge. This phenomena matches the implications of Theorem 1, where $\beta$ must satisfy $\beta\in[\alpha,1]$ to guarantee convergence.
@@ A.2 Effects of Independent Samples’ Batch Size
In this section, we study the effects of batch size for independent samples used in IAN-SGD. For Phase retrieval, we keep $|B|=64$ and other parameters same as section A.1, and we vary independent batch sizes to be $|B^{\prime}|=4,8,16,32,64$. Similarly, for DRO, we keep $|B|=128$ and other parameters same as section A.1, and we vary batch size of independent samples $|B^{\prime}|=16,32,64,128$.
@@
It can be observed that for Phase retrieval and DRO problem, small batch sizes such as $|B^{\prime}|=4,8$ are sufficient to guarantee convergence, whereas large batch sizes of $|B^{\prime}|$ do not significantly increase the sample complexity, which verifies the effectiveness of independent sampling.
@@ A.3 Effects of $\delta$
In Theorem 2, we set $\delta=\frac{\tau_{2}}{1-\tau_{1}}$ based on Assumption 4, which assumes an affine bound for approximation error $\|\nabla f_{\xi}(w)-\nabla F(w)\|$. This assumption relaxes the strong assumption used in Zhang et al. (2019); Liu et al. (2022b); Zhang et al. (2020a), where they assume that the approximation error $\|\nabla f(w)-\nabla F(w)\|$ is upper bounded by a constant. The major weakness is that if certain samples have outlier gradients, such constant upper bound can be very large. Thus, to verify the effectiveness of Assumption 4, we expect convergence of IAN-SGD under a wide range of $\delta$, especially for small $\delta$.
To study the effect of $\delta$, we keep other parameters the same as mentioned in Section A.1 and only vary $\delta$. For Phase Retrieval, we vary $\delta=\{1e^{-8},1e^{-3},1e^{-1},1,10\}$ and Figure 7 (left) shows the corresponding convergence result. For DRO, we vary $\delta=\{1e^{-8},1e^{-3},1e^{-1},1,10\}$ and Figure 7 (right) shows the convergence result. We observe that IAN-SGD convergence is robust to the choice of $\delta$. But $\delta=1e^{-3},1e^{-1}$ demonstrate faster convergence than others for Phase retrieval and DRO, respectively. This result indicates that a small $\delta$ is sufficient to guarantee convergence, which verifies the effectiveness of assumption 4 and IAN-SGD when dealing with stochastic nonconvex generalized-smooth geometry.
@@ A.3.1 Effects of $\tau_{1}$, $\tau_{2}$ on convergence
In Theorem 2, we set $A=\frac{1}{1-\tau_{1}}$ and $\delta=\frac{\tau_{2}}{1-\tau_{1}}$, making the sample complexity dependent on both $\tau_{1}$ and $\tau_{2}$. To further investigate the effect of the gradient noise assumption, we perform experiments that allow flexible choices of $\tau_{1}$ and $\tau_{2}$. Specifically, we control the stochastic gradient noise by computing gradients over mini-batch samples with $B=B^{\prime}=16$ at each iteration. We then evaluate IAN-SGD under various noise settings with $\tau_{1}\in\{0,0.5,0.8\}$ and $\tau_{2}\in\{10^{-8},10^{-3},10^{-1},1,10\}$. As a result, the scaling factor becomes $A\in\{1,2,5\}$ and we define $\delta=A\cdot\{10^{-8},10^{-3},10^{-1},1,10\}$ accordingly. For phase retrieval, we use a unified learning rate $\gamma=1e^{-1}$, and for DRO, we set $\gamma=5e^{-3}$. Other parameters, such as the clipping threshold and $\beta$ keep unchanged. Figure 8 plots the results, where the first row shows IAN-SGD’s convergence for phase retrieval, and the second row shows IAN-SGD’s convergence for DRO.
@@
In this subsection, we study the effects of batch size in IAN-SGD. Specially, we want to justify if IAN-SGD works with $B,B^{\prime}=\Omega(1)$ in practice. For phase retrieval, we set $B=B^{\prime}$ and keep other parameters the same as in section A.1. For comparison, we vary batch sizes to be $\{1,2,4,8,16,32\}$. To guarantee algorithm converging to same level of magnitude, we fine-tune learning rates accordingly, i.e., for each batch size, we set $\gamma=\{8e^{-4},3e^{-2},0.1,0.2,0.9\}$ respectively. Similarly, for DRO, we set $|B|=|B^{\prime}|$ and keep other parameters the same as section A.1. For comparison, we vary batch size to be $\{1,2,4,8,16,32\}$. We fine-tune learning rates accordingly, i.e., $\gamma=\{8e^{-4},3e^{-2},0.1,0.2,0.9\}$. Figure 6 plots the convergence of IAN-SGD under different batch size choices for Phase Retrieval (left) and DRO (right). We conclude IAN-SGD converges under constant-level batch size $B,B^{\prime}=\Omega(1)$.
@@
Next, we compare IAN-SGD with other baseline methods under small batch size. To make a trade-off between convergence and stability, we set $|B|=2$ for SGD, NSGD with momentum, Clip-SGD and SPIDER; For SPIDER algorithm (Fang et al., 2018), we set epoch size to be $20$. For phase retrieval, we used fine-tuned learning rate for all algorithms, i.e., $\gamma=1e^{-5}$ for SGD, $\gamma=1e^{-2}$ for NSGD, NSGD with momentum, $\gamma=5e^{-3}$ for clipped SGD and IAN-SGD, $\gamma=5e^{-2}$ for SPIDER. Similarly, for DRO, we set $|B|=2$ for SGD, NSGD with momentum, Clip-SGD and SPIDER. We used fine-tuned learning rate for all algorithms, i.e., $\gamma=1e^{-6}$ for SGD, $\gamma=1e^{-3}$ for NSGD, NSGD with momentum, clipped SGD and IAN-SGD, $\gamma=5e^{-3}$ for SPIDER. We keep other parameters the same as in Section A.1. The first row of Figure 10 plots the comparison results of Phase Retrieval (left) and DRO (right). We find in both problems, normalized methods such as NSGD, SPIDER, clipped SGD are slow to converge under small batch size.
To further justify the effectiveness of IAN-SGD under small batch size, we also test IAN-SGD and other baseline methods using small batch size to train ResNet He et al. (2016) over CIFAR10 Krizhevsky (2009). For all algorithms, we unify the batch size $B=32$. We use the fine-tuned learning for all algorithms, i.e., $\gamma=1e^{-5}$ for SGD, $\gamma=1e^{-4}$ for Adam, $\gamma=1e^{-3}$ for Adagrad, $\gamma=5e^{-2}$ for NSGD and NSGD with momentum, $\gamma=8e-2$ for clipped SGD and IAN-SGD, the second row of Figure 10 plots the corresponding results for training ResNet18 (left) and ResNet50 (right). As a result, we found IAN-SGD still outperforms than most baselines. This verifies the effectiveness of using IAN-SGD with small batch size in practice.
@@ Appendix B Proof of Descent Lemma 1
where (i) applies Cauchy-schwarz inequality on $\langle\nabla f(w_{\theta})-\nabla f(w),w_{\theta}-w\rangle$, and (ii) is due to Assumption 1, $\|\nabla f(w_{\theta})-\nabla f(w)\big{\|}\leq(L_{0}+L_{1}\|\nabla f(w)\|^{\alpha})\|w_{\theta}-w\|=\theta(L_{0}+L_{1}\|\nabla f(w)\|^{\alpha})\|w^{\prime}-w\|$.
@@ Appendix C Proof for Nonconvex Phase Retrieval
$F(w)=\frac{1}{m}\sum_{\xi\in D}f_{\xi}(w)=\frac{1}{2m}\sum_{\xi\in D}\big{(}y_{\xi}-\big{|}a_{\xi}^{\top}w\big{|}^{2}\big{)}^{2},$ (24)
where $f_{\xi}(w)=\frac{1}{2}(y_{\xi}-|a_{\xi}^{T}w|^{2})^{2}$.
We prove that for every sample $\xi$, $\nabla f_{\xi}(w)=(y_{\xi}-|a_{\xi}^{\top}w|^{2})(a_{\xi}^{\top}{a_{\xi}})w)$ satisfies equation 4 with $\alpha=\frac{2}{3}$. And thus $F(w)$ satisfies expected versions of equation 4, namely
$\mathbb{E}_{\xi}\big{\|}\nabla f_{\xi}(w)-f_{\xi}(w^{\prime})\big{\|}\leq\big{\|}w-w^{\prime}\big{\|}\mathbb{E}_{\xi}(L_{0}+L_{1}\cdot\frac{\|\nabla f_{\xi}(w)\|+\|\nabla f_{\xi}\|^{\alpha}}{2})$ (25)
@@
where (i) extracts scalar $\Big{|}\big{|}a_{\xi}^{\top}w\big{|}^{3}-y_{\xi}\big{|}a_{\xi}^{\top}w\big{|}\Big{|}^{\frac{2}{3}}$ out and use the fact $\frac{1}{2^{{2}/{3}}}\geq\frac{1}{2}$; (ii) uses inequality $|a-b|^{\frac{2}{3}}\geq|a|^{\frac{2}{3}}-|b|^{\frac{2}{3}}$ for any $a,b\in\mathbf{R}$; (iii) applies young’s inequality $|y_{\xi}|^{\frac{2}{3}}|a_{\xi}^{\top}w|^{\frac{2}{3}}\leq\frac{1}{3}\|a_{\xi}^{\top}w\|^{2}+\frac{2}{3}|y_{\xi}|$.
@@
$=$ $\frac{1}{2}\big{\|}2\big{(}\big{|}a_{\xi}^{\top}w^{\prime}\big{|}^{2}-y_{\xi}\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w^{\prime}-2\big{(}\big{|}a_{\xi}^{\top}w\big{|}^{2}-y_{\xi}\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w+\big{(}\big{|}a_{\xi}^{\top}w^{\prime}\big{|}^{2}-y\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w-\big{(}\big{|}a_{\xi}^{\top}w\big{|}^{2}-y\big{)}(a_{\xi}a_{\xi}^{\top})w^{\prime}$
$\quad-\big{(}\big{|}a_{\xi}^{\top}w^{\prime}\big{|}^{2}-y\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w+\big{(}\big{|}a_{\xi}^{\top}w\big{|}^{2}-y\big{)}(a_{\xi}a_{\xi}^{\top})w^{\prime}\big{\|}$
$=$ $\frac{1}{2}\big{\|}\big{(}\big{|}a_{\xi}^{\top}w^{\prime}\big{|}^{2}-y_{\xi}\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w^{\prime}-\big{(}\big{|}a_{\xi}^{\top}w^{\prime}\big{|}^{2}-y\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w+\big{(}\big{|}a_{\xi}^{\top}w\big{|}^{2}-y\big{)}(a_{\xi}a_{\xi}^{\top})w^{\prime}-\big{(}\big{|}a_{\xi}^{\top}w\big{|}^{2}-y_{\xi}\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w$
$\quad+\big{(}\big{|}a_{\xi}^{\top}w^{\prime}\big{|}^{2}-y_{\xi}\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w^{\prime}+\big{(}\big{|}a_{\xi}^{\top}w^{\prime}\big{|}^{2}-y\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w-\big{(}\big{|}a_{\xi}^{\top}w\big{|}^{2}-y\big{)}(a_{\xi}a_{\xi}^{\top})w^{\prime}-\big{(}\big{|}a_{\xi}^{\top}w\big{|}^{2}-y_{\xi}\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w\big{\|}$
@@
Thus, $f_{\xi}(w)=\frac{1}{2}(y_{\xi}-|a_{\xi}^{T}w|^{2})^{2}$ satisfies equation 4 with $\alpha=\frac{2}{3}$, $L_{0}=4y_{\max}a_{\max}^{2}$, and $L_{1}=\frac{9}{2}a^{\frac{4}{3}}_{\max}$. Taking expectation over $\xi$, $F(w)$ satisfies equation 25 with $\alpha=\frac{2}{3}$, $L_{0}=4y_{\max}a_{\max}^{2}$, and $L_{1}=\frac{9}{2}a^{\frac{4}{3}}_{\max}$.

### v3

@@ Appendix A Ablation Study
In order to have a comprehensive understanding of the performance of IAN-SGD in practical problems, we conduct ablation studies regarding on important components of IAN-SGD, including adaptive normalization $\beta$, batch size $|B^{\prime}|$ of independent samples and numerical stabilizer $\delta$.
@@ A.1 Effects of $\beta$
In the first experiment, we choose a unified normalization parameter $\beta=\frac{2}{3}$ for all the normalized methods, i.e., NSGD, NSGD with momentum, clipped SGD, SPIDER and IAN-SGD. To guarantee convergence, we fine-tuned the learning rate accordingly, i.e., $\gamma=0.03$ for NSGD and NSGD with momentum, $\gamma=0.05$ for SPIDER, $\gamma=0.17$ for both clipped SGD and IAN-SGD. To make a fair comparison, we keep other parameters unchanged. Figure 4 (left) shows the comparison of objective value versus sample complexity for the Phase Retrieval problem. It can be observed that, by adjusting $\beta=\frac{2}{3}$, the objective value optimized by all algorithms decreases much faster compared with Figure 2. This indicates that adaptive normalization can accelerate convergence. Moreover, compared with other normalization methods, although IAN-SGD requires additional sampling at each iteration, the training loss still decreases faster than NSGD SGD with momentum and SPIDER.
Similarly, for DRO, we unify the normalization parameter of all the normalized methods, i.e., NSGD, NSGD with momentum, clipped SGD, SPIDER and IAN-SGD with $\beta=\frac{2}{3}$. To guarantee algorithm convergence, we fine-tuned learning rate accordingly. For NSGD, NSGDm and SPIDER, we keep learning rate unchanged; For clipped SGD, and IAN-SGD, we set $\gamma=0.08$. To make a fair comparison, we keep other parameters unchanged. Figure 4 (right) shows the comparison of objective value versus sample complexity. It can be observed that, by setting $\beta=\frac{2}{3}$, the objective value optimized by all normalization methods decreases faster than Figure 2. This verifies the effectiveness of adaptive normalization. Although IAN-SGD requires additional sampling, it converges faster than other normalized methods.
@@
In the second experiment, we vary $\beta$ for IAN-SGD and keep other parameters unchanged. Figure 5 shows the convergence result with different $\beta$ for Phase Retrieval and DRO. It can be observed that decreasing $\beta$ accelerates convergence in general. However, small $\beta$ can make convergence unstable. In the DRO experiment shown in Figure 5 (right), we observed that when $\beta=\frac{3}{5}$, the objective value curve fluctuates a lot. If $\beta$ is smaller than $\frac{3}{5}$, the algorithm fails to converge. This phenomenon matches the implications of Theorem 1, where $\beta$ must satisfy $\beta\in[\alpha,1]$ to guarantee convergence.
@@ A.2 Effects of Independent Samples’ Batch Size
In this section, we study the effects of batch size for independent samples used in IAN-SGD. For Phase retrieval, we keep $|B|=64$ and other parameters same as section A.1, and we vary independent batch sizes to be $|B^{\prime}|=\{4,8,16,32,64\}$. Similarly, for DRO, we keep $|B|=128$ and other parameters same as section A.1, and we vary batch size of independent samples $|B^{\prime}|=\{16,32,64,128\}$.
@@
It can be observed that for Phase retrieval and DRO problem, small batch sizes such as $|B^{\prime}|=\{4,8\}$ are sufficient to guarantee convergence, whereas large batch sizes of $|B^{\prime}|$ do not significantly increase the sample complexity, which verifies the effectiveness of independent sampling.
@@ A.3 Effects of $\delta$
In Theorem 2, we set $\delta=\frac{\tau_{2}}{1-\tau_{1}}$ based on Assumption 4, which assumes an affine bound for approximation error $\|\nabla f_{\xi}(w)-\nabla F(w)\|$. This assumption relaxes the strong assumption used in Zhang et al. (2019); Liu et al. (2022b); Zhang et al. (2020a), where they assume that the approximation error $\|\nabla f_{\xi}(w)-\nabla F(w)\|$ is upper bounded by a constant. The major weakness is that if certain samples have outlier gradients, such constant upper bound can be very large. Thus, to verify the effectiveness of Assumption 4, we expect convergence of IAN-SGD under a wide range of $\delta$, especially for small $\delta$.
To study the effect of $\delta$, we keep other parameters the same as mentioned in Section A.1 and only vary $\delta$. For Phase Retrieval, we vary $\delta=\{1e^{-8},1e^{-3},1e^{-1},1,10\}$ and Figure 7 (left) shows the corresponding convergence result. For DRO, we vary $\delta=\{1e^{-8},1e^{-3},1e^{-1},1,10\}$ and Figure 7 (right) shows the convergence result. We observe that IAN-SGD convergence is robust to the choice of $\delta$. But $\delta=1e^{-3},1e^{-1}$ demonstrate faster convergence than others for Phase retrieval and DRO experiments. This result indicates that a small $\delta$ is sufficient to guarantee convergence, which verifies the effectiveness of assumption 4 and IAN-SGD when dealing with stochastic nonconvex generalized-smooth geometry.
@@ A.3.1 Effects of $\tau_{1}$, $\tau_{2}$ on convergence
In Theorem 2, we set $A=\frac{1}{1-\tau_{1}}$ and $\delta=\frac{\tau_{2}}{1-\tau_{1}}$, making the sample complexity dependent on both $\tau_{1}$ and $\tau_{2}$. To further investigate the effect of the gradient noise assumption, we perform experiments that allow flexible choices of $\tau_{1}$ and $\tau_{2}$. Specifically, we control the stochastic gradient noise by computing gradients over mini-batch samples with $B=B^{\prime}=16$ at each iteration. We then evaluate IAN-SGD under various noise settings with $\tau_{1}\in\{0,0.5,0.8\}$ and $\tau_{2}\in\{10^{-8},10^{-3},10^{-1},1,10\}$. As a result, the scaling factor becomes $A\in\{1,2,5\}$ and we define $\delta=A\cdot\{10^{-8},10^{-3},10^{-1},1,10\}$ accordingly. For phase retrieval, we use a unified learning rate $\gamma=1e^{-1}$, and for DRO, we set $\gamma=5e^{-3}$. Other parameters, such as the clipping threshold and $\beta$, remain unchanged. Figure 8 plots the results, where the first row shows IAN-SGD’s convergence for phase retrieval, and the second row shows IAN-SGD’s convergence for DRO.
@@
In this subsection, we study the effects of batch size in IAN-SGD. Specially, we want to justify if IAN-SGD works with $B,B^{\prime}=\Omega(1)$ in practice. For phase retrieval, we set $B=B^{\prime}$ and keep other parameters the same as in section A.1. For comparison, we vary batch sizes to be $\{1,2,4,8,16,32\}$. To guarantee algorithm converging to same level of magnitude, we fine-tune learning rates accordingly, i.e., for each batch size, we set $\gamma=\{5e^{-4},5e^{-3},2e^{-2},4e^{-2},0.10,0.10\}$, respectively. Similarly, for DRO, we set $|B|=|B^{\prime}|$ and keep other parameters the same as section A.1. For comparison, we vary batch size to be $\{1,2,4,8,16,32\}$, and we fine-tuned learning rates accordingly, i.e., for each batch size, we set $\gamma=\{5e^{-4},1e^{-3},1e^{-3},5e^{-3},5e^{-3},1e^{-2}\}$, respectively. Figure 6 plots the convergence of IAN-SGD under different batch size choices for Phase Retrieval (left) and DRO (right). We conclude IAN-SGD converges under constant-level batch size $B,B^{\prime}=\Omega(1)$.
@@
Next, we compare IAN-SGD with other baseline methods under small batch size. To make a trade-off between convergence and stability, we set $|B|=2$ for SGD, NSGD with momentum, Clip-SGD and SPIDER; For SPIDER algorithm (Fang et al., 2018), we set epoch size to be $20$. For phase retrieval, we fine-tuned learning rate for all algorithms, i.e., $\gamma=1e^{-5}$ for SGD, $\gamma=1e^{-2}$ for NSGD, NSGD with momentum, $\gamma=5e^{-3}$ for clipped SGD and IAN-SGD, $\gamma=5e^{-2}$ for SPIDER. Similarly, for DRO, we set $|B|=2$ for SGD, NSGD with momentum, clipped SGD and SPIDER. We fine-tuned learning rate for all algorithms, i.e., $\gamma=1e^{-6}$ for SGD, $\gamma=1e^{-3}$ for NSGD, NSGD with momentum, clipped SGD and IAN-SGD, $\gamma=5e^{-3}$ for SPIDER. We keep other parameters the same as in Section A.1. The first row of Figure 10 plots the comparison results of Phase Retrieval (left) and DRO (right). We find in both problems, normalized methods such as NSGD, SPIDER, clipped SGD are slow to converge under small batch size.
To further justify the effectiveness of IAN-SGD under small batch size, we also test IAN-SGD and other baseline methods using small batch size to train ResNet (He et al., 2016) over CIFAR10 (Krizhevsky, 2009). For all algorithms, we unify the batch size $B=32$. We fine-tuned learning for all algorithms, i.e., $\gamma=1e^{-5}$ for SGD, $\gamma=1e^{-4}$ for Adam, $\gamma=1e^{-3}$ for Adagrad, $\gamma=5e^{-2}$ for NSGD and NSGD with momentum, $\gamma=8e-2$ for clipped SGD and IAN-SGD, the second row of Figure 10 plots the corresponding results for training ResNet18 (left) and ResNet50 (right). As a result, we found IAN-SGD still outperforms than most baselines. This verifies the effectiveness of using IAN-SGD with small batch size in practice.
@@ Appendix B Proof of Descent Lemma 1
where (i) applies Cauchy-schwarz inequality on $\langle\nabla f(w_{\theta})-\nabla f(w),w_{\theta}-w\rangle$, and (ii) is due to Assumption 1, $\|\nabla f(w_{\theta})-\nabla f(w)\big{\|}\leq(L_{0}+L_{1}\|\nabla f(w)\|^{\alpha})\|w_{\theta}-w\|=\theta(L_{0}+L_{1}\|\nabla f(w)\|^{\alpha})\|w^{\prime}-w\|$. Re-organizing the above inequality gives the desired result.
@@ Appendix C Proof for Nonconvex Phase Retrieval
$F(w)=\frac{2}{m}\sum_{\xi\in D}f_{\xi}(w)=\frac{1}{2m}\sum_{\xi\in D}\big{(}y_{\xi}-\big{|}a_{\xi}^{\top}w\big{|}^{2}\big{)}^{2},$ (24)
where $f_{\xi}(w)=\frac{1}{4}(y_{\xi}-|a_{\xi}^{T}w|^{2})^{2}$.
We prove that for every sample $\xi$, $\nabla f_{\xi}(w)=-(y_{\xi}-|a_{\xi}^{\top}w|^{2})(a_{\xi}{a_{\xi}}^{\top})w$ satisfies equation 4 with $\alpha=\frac{2}{3}$. And thus $F(w)$ satisfies finite-sum versions of equation 4, namely
$\frac{2}{m}\sum_{\xi\in D}\big{\|}f_{\xi}(w)-f_{\xi}(w^{\prime})\big{\|}\leq\big{\|}w-w^{\prime}\big{\|}\cdot\Big{(}L_{0}+\frac{L_{1}}{m}\sum_{\zeta\in D}\frac{\|\nabla f_{\xi}(w)\|^{\alpha}+\|\nabla f_{\xi}(w^{\prime})\|^{\alpha}}{2}\Big{)}$ (25)
@@
where (i) extracts scalar $\Big{|}\big{|}a_{\xi}^{\top}w\big{|}^{3}-y_{\xi}\big{|}a_{\xi}^{\top}w\big{|}\Big{|}^{\frac{2}{3}}$ out; (ii) uses inequality $|a-b|^{\frac{2}{3}}\geq|a|^{\frac{2}{3}}-|b|^{\frac{2}{3}}$ holding for any $a,b\in\mathbf{R}$ (sub-additive property of function $f(x)=x^{2/3}$); (iii) applies young’s inequality $|y_{\xi}|^{\frac{2}{3}}|a_{\xi}^{\top}w|^{\frac{2}{3}}\leq\frac{1}{3}\|a_{\xi}^{\top}w\|^{2}+\frac{2}{3}|y_{\xi}|$.
@@
$=$ $\frac{1}{2}\big{\|}2\big{(}\big{|}a_{\xi}^{\top}w^{\prime}\big{|}^{2}-y_{\xi}\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w^{\prime}-2\big{(}\big{|}a_{\xi}^{\top}w\big{|}^{2}-y_{\xi}\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w+\big{(}\big{|}a_{\xi}^{\top}w^{\prime}\big{|}^{2}-y_{\xi}\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w-\big{(}\big{|}a_{\xi}^{\top}w\big{|}^{2}-y_{\xi}\big{)}(a_{\xi}a_{\xi}^{\top})w^{\prime}$
$\quad-\big{(}\big{|}a_{\xi}^{\top}w^{\prime}\big{|}^{2}-y_{\xi}\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w+\big{(}\big{|}a_{\xi}^{\top}w\big{|}^{2}-y_{\xi}\big{)}(a_{\xi}a_{\xi}^{\top})w^{\prime}\big{\|}$
$=$ $\frac{1}{2}\big{\|}\big{(}\big{|}a_{\xi}^{\top}w^{\prime}\big{|}^{2}-y_{\xi}\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w^{\prime}-\big{(}\big{|}a_{\xi}^{\top}w^{\prime}\big{|}^{2}-y_{\xi}\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w+\big{(}\big{|}a_{\xi}^{\top}w\big{|}^{2}-y_{\xi}\big{)}(a_{\xi}a_{\xi}^{\top})w^{\prime}-\big{(}\big{|}a_{\xi}^{\top}w\big{|}^{2}-y_{\xi}\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w$
$\quad+\big{(}\big{|}a_{\xi}^{\top}w^{\prime}\big{|}^{2}-y_{\xi}\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w^{\prime}+\big{(}\big{|}a_{\xi}^{\top}w^{\prime}\big{|}^{2}-y_{\xi}\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w-\big{(}\big{|}a_{\xi}^{\top}w\big{|}^{2}-y_{\xi}\big{)}(a_{\xi}a_{\xi}^{\top})w^{\prime}-\big{(}\big{|}a_{\xi}^{\top}w\big{|}^{2}-y_{\xi}\big{)}\big{(}a_{\xi}a_{\xi}^{\top}\big{)}w\big{\|}$
@@
Thus, $f_{\xi}(w)=\frac{1}{4}(y_{\xi}-|a_{\xi}^{T}w|^{2})^{2}$ satisfies equation 4 with $\alpha=\frac{2}{3}$, $L_{0}=4y_{\max}a_{\max}^{2}$, and $L_{1}=\frac{9}{2}a^{\frac{4}{3}}_{\max}$. Thus, $F(w)$ satisfies equation 25 with $\alpha=\frac{2}{3}$, $L_{0}=8y_{\max}a_{\max}^{2}$, and $L_{1}={9}a^{\frac{4}{3}}_{\max}$.